OpenStudy (anonymous):

Solve the equation x2 = 49 x = 7

OpenStudy (anonymous):

square root the first \[\sqrt{x^2}=\sqrt{49}\]

OpenStudy (anonymous):

Then drvide?

OpenStudy (anonymous):

no, once you square root it you will have your answer.

OpenStudy (anonymous):

Wait so I square root what exactly? Radicals confuse me.

OpenStudy (anonymous):

whatever is under the radical, radicals mean to square root.

OpenStudy (anonymous):

square root 49

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Haha. Okay thank you. :)

OpenStudy (anonymous):

no prob!